OpenStudy (anonymous):

Organize the following polynomial expressions from least to greatest based on their degree: 6x2 18x3 + 5ab - 6y 8a - 5 4x3y + 3x2 - xy - 4

OpenStudy (lieutenantgeneral):

6x2 = 12 18x3 + 5ab - 6y = 5ab - 6y + 54 8a - 5 = a = 5/8 = 0.625 = 6250 4x3y + 3x2 - xy - 4 = -xy + 12y + 2

OpenStudy (lieutenantgeneral):

u order them

OpenStudy (anonymous):

so would it be this III, I, IV, II ?